A bolt machine produced 250 bolts, of which, 50 are defective. If two bolts are picked at random with replacement, what is the probability that both are defective bolts? *




Express your final answer as a fraction in the lowest term (e.g. 1/4)

1
Expert's answer
2022-03-29T11:59:49-0400

Probability of defective bolt is 50/250=1/5

If we take 2 bolts, probability to get 2 defective bolt is 1/5x1/5=1/25